# AI-Generated Compromises for Coalition Formation: Modeling, Simulation, and a Textual Case Study

### v2

@@ 1.1 Model State
[…]
@@ 1.2 Initialization, Iterative Process, and Halting Conditions
Now, a constitution $const$ is defined as follows.
@@
A constitution $const$ gets as input a tuple $(d_{i},d_{j},p)$ by the mediator and, when applied on $d_{i}$ – and based on the votes of the agents in $d_{i}$, as described by $\{vote(v,p):v\in d_{i}\}$ – returns an assignment to a coalition for each $v\in d_{i}$, namely $const(v)\in\{d_{i},d^{p}\}$, where $d^{p}$ describes the coalition to be possibly-formed around the suggested compromise point $p$.
Following a suggestion of $(d_{i},d_{j},p)$ and an application of the constitution $const$ on $d_{i}$ and $d_{j}$ (which internally depends on the votes of the agents in both coalitions), the resulting Markov state contains a new coalition structure $D^{\prime}$ that is defined as follows: [^2] A coalition with no members can safely be removed from a coalition structure (such that $d^{\prime}_{i}$, $d^{\prime}_{j}$, and $d^{p}$ may be empty). $D^{\prime}:=D\setminus\{d_{i},d_{j}\}\cup\{d^{\prime}_{i},d^{\prime}_{j},d^{p}\},\textrm{where }d^{\prime}_{i}:=(\{v\in d_{i}:const(v)=d_{i}\},p^{i})\textrm{; }d^{\prime}_{j}:=(\{v\in d_{j}:const(v)=d_{j}\},p^{j})\textrm{; }d^{p}:=(\{v\in d_{i}\cup d_{j}:const(v)=d^{p}\},p)$.
@@ 4.1 Modeling Mediators in a Textual Space
Our general solution works as follows. We use word embedding (a standard NLP technique) to translate texts into numerical-valued vectors; this is crucial as, after applying such a word embedding, we are then able to compute distances between the embedded coalition ideal points and use the mediators of the Euclidean space. In this work, we use Google’s Universal Sentence Encoder [7]: this is a pre-trained model that converts sentences into fixed-size vectors, capturing their semantic meanings. (The Universal Sentence Encoder is designed to generate 512-dimensional embedding vectors, providing a semantic representation of sentences.) Thus our embedded metric space contains as elements all those 512512512-length vectors that can be the output of the Universal Sentence Encoder. As a distance measure in this space, we use the square root of the cosine-based dissimilarity [30], a commonly used pseudo metric in NLP. [^6] Formally, we define $d_{\text{cos}}(A,B):=\sqrt{2-2\cdot\frac{A\cdot B}{\|A\|\cdot\|B\|}}\in[0,2],$ which corresponds to $\sqrt{2}\cdot\sin(\theta/2)$, where $\theta$ is the angle between vectors $A$ and $B$. This is a metric when restricted to the unit sphere (or more generally, to rays through the origin), but only a pseudo metric over the entire embedding space: co-linear vectors have distance zero, even if they differ in magnitude. Since our mediator constructs compromise vectors via (non-normalized) weighted averages, we do not assume that vectors lie on a sphere. The mediator guides the coalition formation by proposing sentences to two coalitions within the given word limit (in our simulations, 151515 words). In each iteration, the mediator’s goal is to find two coalitions to suggest a sentence that minimizes the squared cosine similarity pseudo metric between the embedding vector of the chosen sentence and the weighted average of the embedding vectors of the two coalition points.
Our approach to coalition formation in the domain of text relies on the integration of OpenAI’s GPT-3.5-turbo-1106 model (https://openai.com/blog/chatgpt) with a temperature parameter (responsible for the randomness of results) of 0.750.750.75 as recommended in some of the documentation to provide a good trade-off for applications like ours where the output should be coherent but still allow for some diversity and creativity. This LLM takes 333 key roles within our framework:
@@ 4.2 Simulations-Based Analysis
  • Parameter Tests and Scale: The simulations were conducted with different numbers of agents, specifically $n\in\{10,20,30,40,50,100,1000\}$. We varied the parameters $\sigma\in\{0,1,1.5\}$ and $\alpha\in\{-1,0,1\}$, while also setting the boolean variable $C$— enforcing coalition discipline or not. Each parameter combination was tested across 50 repetitions, with all ideal sentences generated sharing a predetermined topic of ways to address global warming.
@@
  • Sentence Selection Process: For each proposed sentence to the two coalitions, the LLM was tasked with generating 101010 sentences that effectively combined both coalition sentences. We followed best practices for structured prompt design and multi-step reasoning [19], including these concepts:
@@
    We used the following prompts and messages given to GPT 3.5 (555 options in total):
    • Mediator 1: Prompt: Generate 101010 possible different well-structured sentences that aggregate the following two sentences. Make sure each sentence has at most 15 words. Number your answers (i.e., 1), 2), 3), 4), 5), and so on) for each sentence you propose. Message: You are a mediator trying to find agreed wording for how to deal with global warming based on existing sentences. Give a straightforward answer with no introduction to help people reach an agreed wording of a coherent sentence. (The proposed sentence was selected based on the minimal squared cosine similarity pseudo metric between its embedding and the weighted average embedding vector, considering the two embedding vectors of the coalitions and their sizes.)
@@
[…]
@@ A.1 A Concrete Example
    Iteration 1: The mediator suggests the compromise point $p^{BC}$ to the coalitions $(C_{B},p^{B})$ and $(C_{C},p^{C})$. Both agents approve since 1<61<61 < 6 and 1<81<81 < 8. We arrive to the following coalition structure $D^{\prime}$:
@@ A.2.1 Simulation-Based Analysis
  • Status Quo: Generated uniformly at random between $\times(0,200)$ ( 0 , 200 ).
  • Ideal Points: Drawing inspiration from the literature [9], each agent was assigned an ideal point $(x,y)$ with both coordinates sampled from either the uniform distribution between (0,200)(0,200)( 0 , 200 ) or from a 2-dimensional Gaussian Mixture Model (GMM). The GMM represents the overall probability distribution as a weighted sum of several Gaussian components with multiple peaks. In our simulations, we considered GMMs with $g$ combined Gaussian distributions, for $g\in\{0,1,2,3,4\}$ with the mean of each Gaussian being distributed uniformly between 0 and 200200200 in each dimension, its deviation distributed uniformly between 0 and 505050, and the weights signifying the importance of each Gaussian are distributed from the $Dirichlet(\alpha^{g}\in{\mathbb{R}_{0}^{+}}^{g})$ distribution with $\alpha$ set to 111 (resulting in $g$ numbers that sum to 111). This sampling of ideal points is demonstrated in the supplementary material. Note that we treat GMM with $g=0$ (i.e., 0 peaks) as the uniform distribution.
@@
We conducted 100100100 independent repetitions for each configuration. Next we present our two evaluation metrics (the first measures the process speed, while the second measures the process quality):
@@
For efficiency, we halt our simulations whenever the number of iterations exceeds a threshold of 10,00010,00010 , 000 (i.e., we treat an instance for which no convergence is reached within 10,00010,00010 , 000 iterations as an instance that does not converge at all).
@@ A.2.2 Results and Discussion
[…]

### v3

@@ 1.1 Model State
    We assume $z\in[n]:=\{1,\ldots,n\}$; that is, the number of coalitions does not exceed the number of agents. The notation $[z]$ refers to the index set of the current coalition structure.
@@ 1.2 Initialization, Iterative Process, and Halting Conditions
Now, a constitution constconst is defined as follows.
@@
A constitution constconst gets as input a tuple $(d_{i},d_{j},p)$ by the mediator and, when applied on $d_{i}$ – and based on the votes of the agents in $d_{i}$, as described by $\{vote(v,p):v\in d_{i}\}$ – returns an assignment to a coalition for each $v\in d_{i}$, namely $const(v)\in\{d_{i},d^{p}\}$, where $d^{p}$ describes the coalition to be possibly-formed around the suggested compromise point $p$.
Following a suggestion of $(d_{i},d_{j},p)$ and an application of the constitution constconst on $d_{i}$ and $d_{j}$ (which internally depends on the votes of the agents in both coalitions), the resulting Markov state contains a new coalition structure $D^{\prime}$ that is defined as follows: [^2] A coalition with no members can safely be removed from a coalition structure (such that $d^{\prime}_{i}$, $d^{\prime}_{j}$, and $d^{p}$ may be empty). $D^{\prime}:=D\setminus\{d_{i},d_{j}\}\cup\{d^{\prime}_{i},d^{\prime}_{j},d^{p}\},\textrm{where }d^{\prime}_{i}:=(\{v\in d_{i}:const(v)=d_{i}\},p^{i})\textrm{; }d^{\prime}_{j}:=(\{v\in d_{j}:const(v)=d_{j}\},p^{j})\textrm{; }d^{p}:=(\{v\in d_{i}\cup d_{j}:const(v)=d^{p}\},p)$.
@@ 4.1 Modeling Mediators in a Textual Space
Our general solution works as follows. We use word embedding (a standard NLP technique) to translate texts into numerical-valued vectors; this is crucial as, after applying such a word embedding, we are then able to compute distances between the embedded coalition ideal points and use the mediators of the Euclidean space. In this work, we use Google’s Universal Sentence Encoder [7]: this is a pre-trained model that converts sentences into fixed-size vectors, capturing their semantic meanings. (The Universal Sentence Encoder is designed to generate 512-dimensional embedding vectors, providing a semantic representation of sentences.) Thus our embedded metric space contains as elements all those $512$-length vectors that can be the output of the Universal Sentence Encoder. As a distance measure in this space, we use the square root of the cosine-based dissimilarity [30], a commonly used pseudo metric in NLP. [^6] Formally, we define $d_{\text{cos}}(A,B):=\sqrt{2-2\cdot\frac{A\cdot B}{\|A\|\cdot\|B\|}}\in[0,2],$ which corresponds to $\sqrt{2}\cdot\sin(\theta/2)$, where $\theta$ is the angle between vectors $A$ and $B$. This is a metric when restricted to the unit sphere (or more generally, to rays through the origin), but only a pseudo metric over the entire embedding space: co-linear vectors have distance zero, even if they differ in magnitude. Since our mediator constructs compromise vectors via (non-normalized) weighted averages, we do not assume that vectors lie on a sphere. The mediator guides the coalition formation by proposing sentences to two coalitions within the given word limit (in our simulations, $15$ words). In each iteration, the mediator’s goal is to find two coalitions to suggest a sentence that minimizes the squared cosine similarity pseudo metric between the embedding vector of the chosen sentence and the weighted average of the embedding vectors of the two coalition points.
Our approach to coalition formation in the domain of text relies on the integration of OpenAI’s GPT-3.5-turbo-1106 model (https://openai.com/blog/chatgpt) with a temperature parameter (responsible for the randomness of results) of $0.75$ as recommended in some of the documentation to provide a good trade-off for applications like ours where the output should be coherent but still allow for some diversity and creativity. This LLM takes $3$ key roles within our framework:
@@ 4.2 Simulations-Based Analysis
  • Parameter Tests and Scale: The simulations were conducted with different numbers of agents, specifically $n\in\{10,20,30,40,50,100,1000\}$. We varied the parameters $\sigma\in\{0,1,1.5\}$ and $\alpha\in\{-1,0,1\}$, while also setting the boolean variable CC— enforcing coalition discipline or not. Each parameter combination was tested across 50 repetitions, with all ideal sentences generated sharing a predetermined topic of ways to address global warming.
@@
  • Sentence Selection Process: For each proposed sentence to the two coalitions, the LLM was tasked with generating $10$ sentences that effectively combined both coalition sentences. We followed best practices for structured prompt design and multi-step reasoning [19], including these concepts:
@@
    We used the following prompts and messages given to GPT 3.5 ($5$ options in total):
    • Mediator 1: Prompt: Generate $10$ possible different well-structured sentences that aggregate the following two sentences. Make sure each sentence has at most 15 words. Number your answers (i.e., 1), 2), 3), 4), 5), and so on) for each sentence you propose. Message: You are a mediator trying to find agreed wording for how to deal with global warming based on existing sentences. Give a straightforward answer with no introduction to help people reach an agreed wording of a coherent sentence. (The proposed sentence was selected based on the minimal squared cosine similarity pseudo metric between its embedding and the weighted average embedding vector, considering the two embedding vectors of the coalitions and their sizes.)
@@
Table 1: A comparison of different mediators-each corresponding to different prompts and LLM-usage strategy. The mean number of iterations until convergence is shown, validated with $95\%$ confidence using ANOVA and Tukey HSD.
@@ 5.2 Future Work
Acknowledgment.
The research was partially funded by the European Union, under the Horizon Europe project PERYCLES (Grant Agreement No. 101094862).
@@ A.1 A Concrete Example
    Iteration 1: The mediator suggests the compromise point $p^{BC}$ to the coalitions $(C_{B},p^{B})$ and $(C_{C},p^{C})$. Both agents approve since $1<6$ and $1<8$. We arrive to the following coalition structure $D^{\prime}$:
@@ A.2.1 Simulation-Based Analysis
  • Status Quo: Generated uniformly at random between $(0,200)\times(0,200)$.
  • Ideal Points: Drawing inspiration from the literature [9], each agent was assigned an ideal point $(x,y)$ with both coordinates sampled from either the uniform distribution between $(0,200)$ or from a 2-dimensional Gaussian Mixture Model (GMM). The GMM represents the overall probability distribution as a weighted sum of several Gaussian components with multiple peaks. In our simulations, we considered GMMs with $g$ combined Gaussian distributions, for $g\in\{0,1,2,3,4\}$ with the mean of each Gaussian being distributed uniformly between 0 and $200$ in each dimension, its deviation distributed uniformly between 0 and $50$, and the weights signifying the importance of each Gaussian are distributed from the $Dirichlet(\alpha^{g}\in{\mathbb{R}_{0}^{+}}^{g})$ distribution with $\alpha$ set to $1$ (resulting in $g$ numbers that sum to $1$). This sampling of ideal points is demonstrated in the supplementary material. Note that we treat GMM with $g=0$ (i.e., 0 peaks) as the uniform distribution.
@@
We conducted $100$ independent repetitions for each configuration. Next we present our two evaluation metrics (the first measures the process speed, while the second measures the process quality):
@@
For efficiency, we halt our simulations whenever the number of iterations exceeds a threshold of $10,000$ (i.e., we treat an instance for which no convergence is reached within $10,000$ iterations as an instance that does not converge at all).
@@ A.2.2 Results and Discussion
Next we discuss the main conclusions, drawn at a $5\%$ significance level: